Hospital waiting room times are normally distributed with a mean of 38.12 minutes and a standard deviation of 8.63 minutes. What is the shortest wait time that would still be in the worst 20% of wait times?
1
Expert's answer
2020-09-29T18:05:50-0400

solution


waiting times follow a normal distribution with


"Mean, \mu=38.12""Standard\ deviation,\sigma=8.63"

Longer waiting times are worse than shorter waiting times. Hence the worst 20% of wait times are wait times on the right tail of the distribution. The inferred level of confidence is 0.80.


The z value corresponding to the right tail probability of 0.2 is

"Z=0.85"

But


"Z = \frac{x-\mu}{\sigma}""x =Z*\sigma +\mu""=0.85 * 8.63 +38.12 =45.4555"

answer:


the shortest wait time that would still be in the worst 20% of wait times is 45.4555 minutes